DIFFERENTIAL GEOMETRY/PDE SEMINAR

Thursday, May 20, 2004 Padelford C-36 1:30 pm A background independent approach to quantum gravity

Abhay Ashtekar

(Center for Gravitational Physics and Geometry, Pennsylvania State University)

A quantum theory of geometry was developed in the mid-nineties to serve as a background independent area for quantum gravity theories. The talk will begin with a summary of this theory and then outline its applications to some of the long standing problems related to black holes and the big-bang.

For more information about this seminar, visit the DG/PDE Seminar Web page (from the Math Department home page, www.math.washington.edu, follow the link Seminars, Colloquia, and Conferences).

The University of Washington is committed to providing access, equal opportunity and reasonable accommodation in its services, programs, activities, education and employment for individuals with disabilities. To request disability accommodation contact the Disability Services Office at least ten days in advance at: 206-543-6450/V, 206-543-6452/TTY, 206-685-7264 (FAX), or dso@u.washington.edu.